### v1

# MdEval: Massively Multilingual Code Debugging

Abstract
Code large language models (LLMs) have made significant progress in code debugging by directly generating the correct code based on the buggy code snippet. Programming benchmarks, typically consisting of buggy code snippet and their associated test cases, are used to assess the debugging capabilities of LLMs. However, many existing benchmarks primarily focus on Python and are often limited in terms of language diversity (e.g., DebugBench and DebugEval). To advance the field of multilingual debugging with LLMs, we propose the first massively multilingual debugging benchmark, which includes 3.6K test samples of 18 programming languages and covers the automated program repair (APR) task, the code review (CR) task, and the bug identification (BI) task. Further, we introduce the debugging instruction corpora MdEval-Instruct by injecting bugs into the correct multilingual queries and solutions (xDebugGen). Further, a multilingual debugger xDebugCoder trained on MdEval-Instruct as a strong baseline specifically to handle the bugs of a wide range of programming languages (e.g. “Missing Mut” in language Rust and “Misused Macro Definition” in language C). Our extensive experiments on MdEval reveal a notable performance gap between open-source models and closed-source LLMs (e.g., GPT and Claude series), highlighting huge room for improvement in multilingual code debugging scenarios [^1] https://mdeval.github.io/.
1 Introduction
Large language models (LLMs) (OpenAI, 2023; Touvron et al., 2023a; Yang et al., 2024a) designed for code, such as CodeLlama (Rozière et al., 2023), DeepSeekCoder (Guo et al., 2024a), and QwenCoder (Hui et al., 2024a), are highly effective at code understanding and generation. These capabilities make them particularly useful for debugging, where deep comprehension of code structure and logic is essential. Automated program repair (APR) (Wen et al., 2024) aims to automatically fix software bugs without human involvement, significantly reducing time and costs in development processes.
LLMs have recently shown considerable potential in this area. For instance, CodeX (Chen et al., 2021) and GPT-4 series (OpenAI, 2023) outperforming previous conventional methods have demonstrated promising results on bug benchmarks like QuixBugs (Lin et al., 2017). The recent work DebugBench (Tian et al., 2024) creates a debugging benchmark including Python, Java, and CPP for LLM evaluation. However, for diverse programming languages in Figure 1, the multilingual debugging scenario poses more language-specific challenges for APR. Multilingual issues (e.g. “Misused Macro Definition” in programming language C, “Missing mut” in Rust, and “Unused Variable” in Go) highlight the complexities and diversities of locating and fixing bugs in the multilingual debugging scenario. Therefore, there is an urgent need to build a truly massively multilingual debugging code benchmark with a wide variety of generic and language-specific bug types.
Figure 1: Massively multilingual evaluation task comprised of three tasks, including code generation, code completion, and code explanation.
To further characterize the debugging performance of LLMs across different programming languages, we introduce MdEval, a framework for data construction, evaluation benchmark, and a multilingual debugging baseline xDebugCoder, to advance code debugging development. First, we propose MdEval, the first massively multilingual evaluation benchmark for code debugging covering 18 programming languages and 3.6K human-curated samples to assess the capabilities of LLMs across a wide range of languages. Further, we create MdEval-Instruct, a multilingual debugging instruction corpus in 18 languages to help the LLM fix the bug given the buggy code snippet. Specifically, we propose xDebugGen to create the buggy and correct code pair for debugging instruction tuning. The bugs are injected into the queries and solutions with our designed three strategies (1) Injecting bugs into query. (2) Injecting bugs into solution. (3) Injecting bugs with the round-trip code translation. Leveraging MdEval-Instruct, we develop xDebugCoder as a strong baseline, assessing the transferability of LLMs in multilingual debugging tasks.
The contributions are summarized as follows: (1) We propose MdEval, a comprehensive multilingual code debugging benchmark consisting of 3.6K samples spanning three tasks: automated program repair (APR), code review (CR), and bug identification (BI). This benchmark covers 18 languages and includes both generic and language-specific bug types. (2) We introduce the massively multilingual code debugging instruction corpora MdEval-Instruct created by the xDebugGen. By injecting bugs into the multilingual correct query or response, we can create the pairs of buggy code and the correct code for instruction. (3) We systematically evaluate the multilingual code debugging capabilities of 15+ models on our created MdEval and create a leaderboard to evaluate them on 18 programming languages dynamically. Notably, extensive experiments suggest that comprehensive multilingual multitask evaluation can realistically measure the gap between open-source (e.g. DeepSeekCoder and Qwen-Coder) and closed-source models (e.g. GPT-4o and Claude series).
2 Multilingual Code Debugging Evaluation: MdEval
Table 1: MdEval dataset statistics.
Statistics Number
Problems $3,513$
Automated Program Repair $1,171$
Code Review $1,171$
Bug Identification $1,171$
Total Test Cases $6,838$
#Difficulty Level
- Easy/Medium/Hard $1,692$/$1,209$/$612$
Length
Question
      - maximum length $291$ tokens
      - minimum length $7$ tokens
      - avg length $83.3$ tokens
Buggy code
      - maximum length $5913$ tokens
      - minimum length $15$ tokens
      - avg length $239.4$ tokens
2.1 Data Overview
As illustrated in Table 1, the MdEval consists of 3.6K problems. Following Yang et al. (2024c), we design 3 multilingual debugging-related tasks: Automated Program Repair, Code Review, and Bug Identification. Each task contains about 1.2K questions, with more than 60 problems in each language. Each problem in MdEval includes question, example test cases, buggy code, correct code, and unit tests.
We calculated the token lengths of the question and buggy code using the CodeLlama tokenizer (Rozière et al., 2023). The average question length is 83 words, highlighting their detailed descriptive nature. The average buggy code length is 239 tokens, indicating the complexity of the code. In addition, the total number of unit tests for the dataset is 6,838, to ensure the accuracy of the bug-fix judgment. Figure 3 plots error types distribution. We strive to cover all error types in each language. Due to the inherent differences among languages, we ensure a balanced distribution of difficulty levels, leading to variations in the distribution of error types across languages.
(a) Generic Error Types
(b) Language-Specific Error Types
Figure 2: Error types in MdEval. Part (a) shows generic error types, and Part (b) lists language-specific error types.
In Table 2, We compare MdEval with other code debugging benchmarks. Our benchmark provides a valuable enhancement to existing ones, significantly expanding the variety of programming languages and introducing language-specific error types, along with a greater number of questions and diverse bug-fixing tasks.
Table 2: Comparison between MdEval and other code debugging benchmarks. MdEval significantly provides a comprehensive multilingual view by expanding the variety of programming languages and language-specific error types.
Benchmark #Languages #Task Size (Easy/Middle/Hard) #Error Types Source of Bugs Language-specific Bugs
DeepFix (Yasunaga and Liang, 2021) 1 1 6,971 4 Collection
Github-Python (Yasunaga and Liang, 2021) 1 1 15K 14 Collection
Bug2Fix (Lu et al., 2021) 1 1 5,835 - Collection
FixEval (Haque et al., 2023) 2 1 43K/243K - Collection
CodeError (Wang et al., 2023) 1 1 4,463 6 Collection
CodeEditorBench (Guo et al., 2024b) 3 1 676/515/716 14 GPT-4 Generation
DebugBench (Tian et al., 2024) 3 1 1,438/1,401/1,414 18 GPT-4 Generation
DebugEval (Yang et al., 2024c) 3 4 1,933/1,903/1,876 18 Collection & GPT-4 Generation
MdEval (Ours) 18 3 1,692/1,209/612 39 Human Annotation
2.2 MdEval Construction & Quality Control
To curate the massively multilingual code debugging evaluation benchmark MdEval, we employ a comprehensive and systematic human annotation process for multilingual code samples. This process is guided by meticulously defined guidelines to guarantee accuracy and consistency.
We initially recruited 13 computer science graduates as multilingual debugging annotators, all proficient in their respective programming languages. After completing a comprehensive training course on annotation methods, the annotators were tasked with defining problems, providing corresponding solutions, and buggy code. Annotators adhered to the following principles: (1) Write a clear problem question and design test cases to ensure that bugs can be effectively identified; (2) Categorize bugs into multiple difficulty levels (easy/medium/hard) based on the complexity of fixing these code.
Figure 4 illustrates the overall process of dataset construction. We began by collecting code snippets from GitHub, which were then extracted and filtered following StarCoder (Li et al., 2023). Prior to the annotation phase, we summarized generic error types and language-specific error types. The three task definitions and corresponding annotation methods were explained in detail. The annotators proceeded to annotate the code according to the identified error types and specified annotation methods. To ensure annotation quality, they evaluated the annotated code based on four criteria: problem difficulty, ambiguity, error type, and solvable. Furthermore, after completing their annotations, each annotator exchanged data with another annotator for cross-refining, aiming to minimize subjective bias and errors. Any discrepancies between annotators were resolved through consensus or with input from senior annotators. Finally, we engaged three volunteers to assess the accuracy of the benchmark (targeting > 90%) and correct errors.
Figure 3: Error types distribution in 18 programming Languages from the MdEval.
Figure 4: Overview of the MdEval construction process. We collected and filtered code snippets from GitHub. Before annotation, we summarized error types. Annotators then labeled the code based on these types. To ensure quality, they used GPT-4o to evaluate the annotations on four criteria: difficulty, ambiguity, error type, and solvable. Finally, they exchanged data with each other to minimize bias and errors.
Figure 5: Examples of multilingual automated program repair (APR), code review (CR), and code identification (CI).
2.3 Instruction Corpora for Code Debugging
To create the instruction corpora, we need to create the pair of the correct code snippet and the buggy code. First, we select the proper code snippet from 18 languages and prompt the code LLM to generate a new question $q^{L_{k}}$ of programming language $L_{k}$. Then, we use the LLM to generate the correct code $c^{L_{k}}$ and filter the low-quality response with an LLM filter and the generated test cases. Therefore, we can regard the $(q^{L_{k}},c^{L_{k}})$ as the correct sample by ensuring the correctness of $c^{L_{k}}$ as much as possible. We propose xDebugGen comprised of the following three strategies to create the code debugging instruction corpora MdEval-Instruct to obtain the fine-tuned LLM xDebugCoder.
Injecting Bugs into Query.
We can prompt the LLM to modify the original question to another similar question with minor differences, where the similar question $q^{\prime L_{k}}$ is further used to generate the answer $\mathcal{M}_{w}(q^{\prime L_{k}})$ by a weak LLM $\mathcal{M}_{w}$ with small size (e.g. Qwen2.5-1.5B). Since there exist differences between the original question $q^{L_{k}}$ and the modified question $q^{\prime L_{k}}$, $(\mathcal{M}_{w}(q^{\prime L_{k}}),c^{L_{k}})$ can be fed into the LLM as source input and target prediction.
Injecting Bugs into Solution.
Another more intuitive method is to directly inject the bugs into the correct code $c^{L_{k}}$. Given the bug type and the correct code snippet, we prompt the LLM to generate the buggy code $\mathcal{M}(c^{L_{k}})$. The pair $(\mathcal{M}(c^{L_{k}}),c^{L_{k}})$ can be used for the instruction tuning.
Injecting Bugs with Round-trip Code Translation.
Under the multilingual scenario, we can translate the correct $c^{L_{k}}$ into the $\mathcal{M}_{w}(c^{L_{k}};L_{k}\to L_{j})$ and then back-translate into the original language $L_{k}$ of programming languages using the weak LLM $\mathcal{M}$, where the round-trip translation code snippet can be regarded as the buggy code. The pair $(\mathcal{M}_{w}(\mathcal{M}_{w}(c^{L_{k}};L_{k}\to L_{j});L_{j}\to L_{k}),c^{L_{k}})$ can be used for the instruction tuning.
2.4 Evaluation Task
Automated Program Repair (APR).
The automated program repair task forces the LLM to fix the bug in the given code snippet and then generates the correct code. Given the programming language $L_{k}\in\{L_{i}\}_{i=1}^{K}$ ($K=18$ is the number of programming languages), we provide the question $q^{L_{k}}$, the corresponding buggy code $b^{L_{k}}$, and the examples test cases $e^{L_{k}}$ for inputs. We can organize the different input settings for evaluation:
$r^{L_{k}}=\mathbb{I}(P(c^{L_{k}}|I;\mathcal{M});u^{L_{k}})$ (1)
where $\mathbb{I}(\cdot)$ is the executor of the multilingual sandbox to verify the correctness of the generated code with the test cases $u^{L_{k}}$ (If the fixed code $c^{L_{k}}$ passes all test cases, the evaluation result $r^{L_{k}}=1$, else 00). In our work, we provide three settings for evaluation to simulate the realistic user queries: (1) Question with buggy code: $I=\{q^{L_{k}},b^{L_{k}}\}$ (2) Buggy code with example test cases: $I=\{b^{L_{k}},e^{L_{k}}\}$ (3) Only buggy code: $I=\{b^{L_{k}}\}$.
Code Review (CR).
For the code review task, we present two versions of code to LLMs: the correct code $b^{L_{k}}$ and the buggy code $c^{L_{k}}$ with only a few minor differences between them. The correct code and buggy code are listed in a random order to feed into LLM for distinguishing the buggy code.
Bug Identification (BI).
In this task, LLMs are required to classify the type of error present in a given buggy program $c^{L_{k}}$ with one error. The LLMs must choose the correct error category from 39 bug types (including generic bug types and language-specific bug types).
2.5 Evaluation Metrics
Automated Program Repair.
In the automated program repair task, we evaluate models by executing the generated code against a set of unit tests and assessing performance using the Pass@1 metric (pass rate for just one-time generation). Greedy Pass@1 indicates whether a result produced by the LLM successfully passes all corresponding unit tests.
Code Review & Bug Identification.
In the code review and bug identification tasks, we evaluate model performance using accuracy, as both require the model to select from a set of provided options.
3 Experiments
3.1 Experiment Setup
Code LLMs.
We evaluate 16 popular models, both closed-source and open-source (sizes ranging from 1.3B to 236B parameters). For general models, we evaluate GPTs (OpenAI, 2023) (GPT4-o, GPT4-o-mini), Claude-3.5 (Anthropic, 2023). For code models, we test Qwen2.5-Coder (Hui et al., 2024b), DeepSeekCoder (DS-Coder) (Guo et al., 2024a), CodeLlama (Rozière et al., 2023), and Codegemma (Gemma Team, 2024). Furthermore, we fine-tune the Qwen2.5-Coder-7B to provide a baseline model xDebugCoder for reference. For closed-source models, the responses are generated by the official API. For the open-source models, we perform inference on all models using the vLLM (Kwon et al., 2023) framework. All models adopt a greedy decoding strategy during inference, the temperature is set to 0, and the maximum generation length is 2048.
xDebugCoder Training Corpora.
The training corpora consist of our debugging dataset MdEval-Instruct, which contains 16K samples, and the Magicoder-Instruct code generation dataset (Wei et al., 2023), comprising 180K samples. This combination ensures that the model possesses a fundamental capability to follow instructions for basic code tasks. We apply data decontamination before training our xDebugGen. Following Li et al. (2023); Wei et al. (2023), we adopt the N-gram exact match decontamination method with MdEval, HumanEval(Chen et al., 2021), MultiPL-E(Cassano et al., 2023), MBPP(Austin et al., 2021).
xDebugCoder Optimization.
Our model, xDebugCoder, based on Qwen2.5-Coder-7B, is trained with full parameters for 3 epochs using a cosine scheduler, starting at a learning rate of $5\times 10^{-5}$ with 3% of total training steps for warmup. We utilize AdamW (Loshchilov and Hutter, 2017) as the optimizer; the batch size is set to 1024, with a maximum sequence length of 2048. All experiments are performed with 8 NVIDIA A800-80GB GPUs.
Table 3: Pass@1 (%) scores of different models for Automated Program Repair tasks on MdEval. The underlined numbers are the best scores for each language. “Avgall” represents the average scores of all code languages.
Model Size Avgall C C# CLISP CPP F# Go JS Java
Closed-Source Models & 200B+ Models
[…]
DS-Coder-V2-Instruct 236B 72.7 60.9 87.1 70.0 63.6 80.0 44.6 61.4 78.3
1B+ Models
DS-Coder-Instrcut 1.3B 33.5 43.8 21.4 16.7 37.9 13.8 22.9 45.7 31.7
Yi-Coder-Chat 1.5B 33.0 35.9 38.6 8.3 39.4 9.2 26.5 45.7 36.7
Qwen2.5-Coder-Instruct 1.5B 36.2 29.7 2.9 13.3 33.3 16.9 10.8 58.6 35.0
6B+ Models
DS-Coder-Instruct 6.7B 52.9 59.4 7.1 53.3 63.6 23.1 37.3 67.1 60.0
CodeLlama-Instruct 7B 31.9 32.8 2.9 23.3 33.3 12.3 15.7 62.9 20
CodeGemma-Instruct 7B 48.9 45.3 2.9 3.3 36.4 20 41 71.4 66.7
CodeQwen1.5-Chat 7B 50.3 37.5 31.4 41.7 43.9 41.5 33.7 55.7 58.3
Qwen2.5-Coder-Instruct 7B 63.4 59.4 57.1 60.0 62.1 29.2 44.6 60.0 85.0
Yi-Coder-Chat 9B 48.3 54.7 7.1 16.7 47.0 16.9 4.8 68.6 80.0
DS-Coder-V2-Lite-Instruct 16B 59.5 60.9 12.9 40.0 57.6 32.3 32.5 65.7 76.7
DS-Coder-Instruct 33B 55.6 70.3 10 56.7 59.1 27.7 14.5 70 76.7
xDebugGen (Our Method) 7B 46.8 53.1 2.9 36.7 54.5 35.4 3.6 54.3 48.3
Method Json Julia PHP Pascal Python R Ruby Rust Scala Swift
Closed-Source Models & 200B+ Models
GPT-4o-240806 20.3 82.1 87.5 72.3 90.2 86.7 84.1 80.0 92.2 78.3
GPT-4o-mini-240718- 39.0 85.1 85.9 64.6 88.5 80.0 84.1 71.4 76.6 70.0
Claude-3.5-sonnet-240620 44.1 82.1 78.1 58.5 90.2 85.0 79.4 91.4 92.2 75.0
Claude-3.5-sonnet-241022 40.7 82.1 85.9 60.0 93.4 86.7 82.5 85.7 89.1 78.3
DS-Coder-V2-Instruct 44.1 83.6 71.9 67.7 85.2 75.0 81.0 85.7 87.5 86.7
1B+ Models
DS-Coder-Instrcut 8.5 46.3 31.2 40.0 49.2 20.0 57.1 32.9 31.2 53.3
Yi-Coder-Chat 16.9 40.3 28.1 24.6 54.1 20.0 52.4 42.9 34.4 36.7
Qwen2.5-Coder-Instruct 28.8 61.2 43.8 46.2 57.4 26.7 57.1 61.4 23.4 50.0
6B+ Models
DS-Coder-Instruct 23.7 67.2 57.8 58.5 78.7 48.3 82.5 71.4 35.9 63.3
CodeLlama-Instruct 11.9 43.3 34.4 38.5 39.3 21.7 63.5 58.6 23.4 36.7
CodeGemma-Instruct 28.9 68.7 48.4 50.8 73.8 46.7 74.6 75.7 64.1 61.7
CodeQwen1.5-Chat 23.7 64.2 53.1 52.3 68.9 45.0 71.4 77.1 37.5 71.7
Qwen2.5-Coder-Instruct 35.6 74.6 75.0 66.2 78.7 65.0 69.8 77.1 76.6 71.7
Yi-Coder-Chat 25.4 65.7 54.7 52.3 67.2 36.7 74.6 68.6 70.3 70.0
DS-Coder-V2-Lite-Instruct 30.5 70.1 68.8 64.6 73.8 71.7 77.8 82.9 79.7 83.3
DS-Coder-Instruct 28.9 70.1 54.7 50.8 73.8 60 85.7 67.1 76.6 63.3
xDebugGen (Our Method) 23.7 61.2 64.1 46.2 63.9 36.7 65.1 74.3 62.5 68.3
Table 4: Accuracy of different models for Code Review tasks on MdEval. The underlined numbers are the best scores for each language. “Avgall” represents the average accuracy of all code languages.
Model Size Avgall C C# CLISP CPP F# Go JS Java
Closed-Source Models
[…]
1B+ Models
DS-Coder-Instrcut 1.3B 54.5 68.8 57.1 51.7 49.2 62.1 45.8 58.6 53.3
Yi-Coder-Chat 1.5B 56.2 50.0 47.1 50.0 50.8 39.4 55.4 81.4 56.7
Qwen2.5-Coder-Instruct 1.5B 51.4 53.1 51.4 48.3 49.2 51.5 53.0 51.4 55.0
6B+ Models
DS-Coder-Instruct 6.7B 48.3 45.3 47.1 46.7 50.8 48.5 48.2 50.0 45.0
CodeQwen1.5-Chat 7B 65.3 68.8 75.7 51.7 63.1 68.2 77.1 84.3 68.3
Qwen2.5-Coder-Instruct 7B 80.0 67.2 90.0 68.3 83.1 74.2 73.5 87.1 85.0
Yi-Coder-Chat 9B 67.1 59.4 70.0 58.3 67.7 63.6 56.6 80.0 70.0
DS-Coder-V2-Lite-Instruct 16B 63.7 70.3 65.7 56.7 60.0 66.7 56.6 71.4 61.7
xDebugGen (Our Method) 7B 60.4 59.4 61.4 68.3 56.1 67.7 65.1 65.7 61.7
Method Json Julia PHP Pascal Python R Ruby Rust Scala Swift
Closed-Source Models
GPT-4o-240806 91.5 91.0 92.2 93.8 98.4 98.3 90.5 94.3 95.3 95.0
GPT-4o-mini-240718 78.0 92.5 89.1 95.4 93.4 95.0 90.5 90.0 82.8 86.7
Claude-3.5-sonnet-240620 89.8 94.0 95.3 95.4 95.1 95.0 95.2 97.1 90.6 96.7
Claude-3.5-sonnet-241022 89.8 92.5 95.3 95.4 98.4 98.3 93.7 100.0 95.3 98.3
1B+ Models
DS-Coder-Instrcut 50.8 52.2 51.6 56.9 49.2 56.7 63.5 55.7 42.2 55.0
Yi-Coder-Chat 62.7 58.2 56.2 70.8 73.8 53.3 41.3 71.4 42.2 50.0
Qwen2.5-Coder-Instruct 52.5 56.7 50.0 47.7 44.3 46.7 57.1 52.9 51.6 53.3
6B+ Models
DS-Coder-Instruct 42.4 44.8 51.6 50.8 52.5 55.0 42.9 50.0 45.3 51.7
CodeQwen1.5-Chat 59.3 53.7 60.9 43.1 72.1 65.0 77.8 62.9 60.9 63.3
Qwen2.5-Coder-Instruct 61.0 85.1 95.3 76.9 73.8 88.3 81.0 74.3 79.7 96.7
Yi-Coder-Chat 52.5 74.6 84.4 63.1 68.9 53.3 69.8 75.7 64.1 75.0
DS-Coder-V2-Lite-Instruct 54.2 61.2 62.5 76.9 62.3 55.0 69.8 68.6 57.8 68.3
xDebugGen (Our Method) 55.9 62.7 60.9 52.3 52.5 60.0 60.3 55.7 54.7 66.7
Table 5: Accuracy of different models for Bug Identification tasks on MdEval. The underlined numbers are the best scores for each language. “Avgall” represents the average accuracy of all code languages.
Model Size Avgall C C# CLISP CPP F# Go JS Java
Closed-Source Models
[…]
1B+ Models
DS-Coder-Instrcut 1.3B 18.4 14.1 20.0 20.0 18.2 9.2 8.4 30.0 23.3
Yi-Coder-Chat 1.5B 1.4 3.1 2.9 3.3 3.0 0.0 0.0 0.0 0.0
Qwen2.5-Coder-Instruct 1.5B 1.2 3.1 2.9 3.3 3.0 1.5 0.0 0.0 0.0
6B+ Models
DS-Coder-Instruct 6.7B 9.3 10.9 2.9 6.7 12.1 0.0 6.0 28.6 8.3
CodeQwen1.5-Chat 7B 5.7 4.7 11.4 10.0 4.5 1.5 1.2 8.6 8.3
Qwen2.5-Coder-Instruct 7B 8.5 12.5 8.6 6.7 12.1 0.0 12.0 11.4 6.7
Yi-Coder-Chat 9B 14.5 25.0 14.3 11.7 18.2 3.1 9.6 18.6 18.3
DS-Coder-V2-Lite-Instruct 16B 13.6 14.1 14.3 8.3 7.6 18.5 14.5 11.4 5.0
xDebugGen (Our Method) 7B 4.6 7.8 8.6 6.7 7.6 0.0 2.4 4.3 1.7
Method Json Julia PHP Pascal Python R Ruby Rust Scala Swift
Closed-Source Models
GPT-4o-240806 18.6 20.9 25.0 30.8 45.9 35.0 46.0 41.4 18.8 13.3
GPT-4o-mini-240718 18.5 11.9 20.3 16.9 32.8 31.7 30.2 22.9 10.9 23.3
Claude-3.5-sonnet-240620 24.6 35.8 31.2 40.0 54.1 48.3 41.3 41.4 26.6 28.3
Claude-3.5-sonnet-241022 23.1 43.3 25.0 35.4 54.1 46.7 38.1 45.7 29.7 23.3
1B+ Models
DS-Coder-Instrcut 10.2 20.9 31.2 9.2 34.4 15.0 33.3 7.1 12.5 15.0
Yi-Coder-Chat 5.1 0.0 0.0 1.5 3.3 0.0 0.0 2.9 0.0 0.0
Qwen2.5-Coder-Instruct 1.7 0.0 0.0 0.0 3.3 0.0 0.0 2.9 0.0 0.0
6B+ Models
DS-Coder-Instruct 6.8 7.5 4.7 1.5 19.7 5.0 25.4 11.4 7.8 1.7
CodeQwen1.5-Chat 6.8 1.5 0.0 6.2 9.8 1.7 14.3 10.0 1.6 0.0
Qwen2.5-Coder-Instruct 15.3 3.0 1.6 6.2 24.6 6.7 12.7 7.1 1.6 3.3
Yi-Coder-Chat 6.8 19.4 6.2 6.2 36.1 8.3 17.5 15.7 14.1 11.7
DS-Coder-V2-Lite-Instruct 3.4 17.9 12.5 12.3 16.4 25.0 15.9 12.9 15.6 20.0
xDebugGen (Our Method) 5.1 3 0.0 1.5 14.8 3.3 9.5 5.7 0.0 0.0
3.2 Main Results
Automated Program Repair.
Table 3 presents the Pass@1 results of different models on MdEval for the multilingual automated program repair task (given question with buggy code, request the model to fix buggy code). The results indicate a marked disparity between closed-source state-of-the-art models and the majority of open-source models across nearly all programming languages. Notably, GPT-4o, Claude-3.5-sonnet, and DS-Coder-V2-236B excel in this task and demonstrate significant performance advantages over other models. Furthermore, our baseline model xDebugCoder, was fine-tuned using only 16K bug-related data MdEval-Instruct. Despite the limited size of this dataset, the model demonstrated competitive performance compared to others of similar scale, highlighting the effectiveness of MdEval-Instruct in enhancing the debugging capabilities of models.
Code Review.
Table 4 displays the accuracy for multilingual code review tasks. The results show that closed-source models still significantly outperform open-source models in the code review task. This task requires the model to distinguish between buggy and correct code. The closed-source models demonstrate a strong ability to understand complex code logic, achieving an accuracy rate of approximately 90%. In contrast, the smaller open-source model exhibits significant challenges, with an accuracy rate of around 50%. This disparity underscores the limitations of the current open-source options in effectively interpreting intricate coding patterns.
Bug Identification.
In the bug identification task, the goal is to identify the error type in a given code snippet, where the LLMs analyze source code for defects and choose the correct bug type from the pre-defined 39 bug types. Table 5 lists the all results of the bug identification. Notably, the closed-source LLMs, such as GPT-4o and Claude series, have the dominant advantages, outperforming the open-source LLMs by nearly +10 points. Bug identification with 39 bug types poses a daunting challenge to the LLMs, requiring alignment capability of LLMs between the given code snippet and its corresponding bug type.
4 Further Analysis
4.1 Performance across error types.
As shown in Figure 6, The performance of models on the automated program repair task varies across different error types, highlighting the strengths and weaknesses of these models in addressing specific challenges. Consistently, the models demonstrate robust capabilities in repairing syntax errors, reference errors, and logic errors. These error types tend to be more straightforward and well-defined, allowing the models to leverage their knowledge effectively to identify and correct issues with high accuracy. In contrast, the models exhibit their worst performance when dealing with language-specific errors. Language-specific errors can arise from unique syntax rules, idiomatic expressions, or even cultural programming practices that are not universally applicable. As a result, addressing these types of errors presents a significant challenge and underscores the need for further improvements in the models’ training processes and methodologies.
4.2 Another Two Automated Program Repair Settings
Figure 6: The performance of models on the automated program repair task across error types.
In Figure 7, we explore two additional automated program repair settings that aim to simulate realistic user queries in software debugging. Part (a) presents the results for the scenario in which models are given both buggy code and corresponding example test cases. This setup allows for a comprehensive evaluation of the ability of models to understand and correct specific issues based on contextual examples. In contrast, Part (b) illustrates the results for a more challenging scenario where only the buggy code is provided to the models, requiring them to identify and rectify errors without any additional context. This comparison highlights the varying capabilities of models in different settings, emphasizing the importance of context in automated program repair.
(a) Buggy code with example test cases
(b) Only buggy code
Figure 7: Two additional automated program repair settings are designed to simulate realistic user queries. Part (a) presents results for the scenario where models are provided with buggy code along with example test cases, while Part (b) illustrates results for the scenario where only the buggy code is provided to the models.
5 Related Work
Code Large Language Model.
With the rapid advancement of large language models(LLMs) (OpenAI, 2023; Touvron et al., 2023b; AI, 2024; Bai et al., 2023; Yang et al., 2024a), solving complex code-related tasks has become increasingly feasible, leading to the emergence of numerous Code LLMs. Early studies utilized models like BERT (Devlin et al., 2019) or GPT (Radford et al., 2018) as backbones, trained on billions of code snippets to enable tasks involving code understanding and generation (Chen et al., 2021; Feng et al., 2020; Scao et al., 2022; Li et al., 2022; Wang et al., 2021; Allal et al., 2023). Recently, advancements in domain-specific pre-training and instruction fine-tuning techniques (Zheng et al., 2024a; Yue et al., 2024) have led to extensive efforts in fine-tuning models on large-scale code corpora and crafting code-related task instructions (Rozière et al., 2023; Zheng et al., 2023; Luo et al., 2023; Muennighoff et al., 2023; Gemma Team, 2024; Zheng et al., 2024b; Guo et al., 2024a; Wei et al., 2023; Sun et al., 2024; Lozhkov et al., 2024; Jiang et al., 2023; Hui et al., 2024b; Wang et al., 2024a; Deng et al., 2024; Liu et al., 2024). These models demonstrate remarkable performance in tasks like code completion, synthesis, and program repair.
Debugging with Large Language Models.
Automatic program debugging holds substantial practical value. With the emergence of LLM capabilities, a growing number of individuals are utilizing LLMs for code debugging, leading to extensive research in this field. Code Debugging includes serval tasks such as bug or vulnerability detection (Pradel and Sen, 2018; Allamanis et al., 2021; Yuan et al., 2023; Zhang et al., 2024; Zhong et al., 2024), fuzz test (Deng et al., 2023; Xia et al., 2024; Yang et al., 2024b), program repair  (Wen et al., 2024; Lin et al., 2017; Zhang et al., 2023a; Prenner and Robbes, 2023; Gu et al., 2024; Tambon et al., 2024; Wang et al., 2024b), GitHub issues auto resolving  (Jimenez et al., 2023; Chen et al., 2024; Tao et al., 2024). To effectively assess the code debugging capabilities of LLMs, several benchmark tests have been introduced  (Prenner et al., 2022; Sobania et al., 2023; Xia and Zhang, 2023; Zhang et al., 2023b; Tian et al., 2024; Yang et al., 2024c). Notably, DebugBench (Tian et al., 2024) provides a comprehensive classification of error types and analyzes the debugging capabilities of LLMs based on these categories. Similarly, DebugEval (Yang et al., 2024c) has designed various debugging-related tasks to evaluate LLM performance across different task dimensions. However, these studies focus on 1 to 3 languages. In reality, there are significant differences in code errors between languages, leading to numerous language-specific errors. To address this gap, we propose MdEval, a comprehensive code debugging benchmark covering 18 languages, aiming to assess LLM debugging capabilities from a broader perspective.
6 Conclusion
In this work, we introduce MdEval of instruction corpora MdEval-Instruct, evaluation benchmark, and a strong baseline xDebugCoder, where the benchmark includes automated program repair (APR), code review (CR), and bug identification (BI) of 18 programming languages (total 3.6K samples), aiming to assess the debugging capabilities of large language models (LLMs) in multilingual environments. Further, we propose xDebugGen to construct a multilingual debugging instruction corpus, where we inject the bugs into the query or answer to create the pair of the buggy code and correct code. Based on MdEval-Instruct, we develop xDebugCoder, a multilingual LLM for debugging in a wide range of programming languages as a strong baseline. Through extensive experiments, this paper reveals a substantial performance gap between open-source and closed-source LLMs, underscoring the need for further improvements in multilingual code debugging. In the future, we will continue expanding the number of languages in MdEval.